### v1

# Seeing and Knowing in the Wild: Open-domain Visual Entity Recognition with Large-scale Knowledge Graphs via Contrastive Learning

Abstract
[…]
Introduction
The ability to recognize and identify visual entities in the open world is a critical milestone for scalable computer vision systems. In contrast to traditional classification tasks that depend on a fixed set of categories, open-world visual entity recognition seeks to recognize and link images to a vast and evolving universe of real-world entities, such as specific landmarks, artworks, biological species, or public figures. The recently introduced Open-domain Visual Entity recognitioN (OVEN) benchmark (Hu et al. 2023) challenges models to link images and accompanying text query (specifying the intent of the image) to the correct Wikidata entities,  [^1] We consistently use the term entity to refer to uniquely identifiable real-world concepts in Wikidata (e.g. Albert Einstein (Q937) and Golden Retriever (Q38686)), which differs from the typical usage of class, denoting categories within a predefined set. identified by its unique QID. A few examples of the OVEN-based tasks are shown in Figure 1a.
Previous approaches addressing open-world visual entity recognition typically adopt a dual-encoder paradigm, which aligns visual representations with textual descriptions of corresponding entities. CLIP2CLIP and CLIPFusion (Hu et al. 2023) include the information of the lead image(s) of the entities into the training pipeline for a better alignment of different modalities. Recently, researchers (Hu et al. 2023; Caron et al. 2024; Xiao et al. 2024) have developed two-step approaches which leverage generative language models, such as PaLI (Chen et al. 2023), GIT (Wang et al. 2022), and Vicuna (Zheng et al. 2023), to generate textual labels of images and then utilize search algorithms (e.g. BM25 (Robertson, Walker, and Hancock-Beaulieu 1995)) to identify entities whose names closely match these predicted labels.
Despite the progress, this task remains fundamentally challenging. First, the label space is extremely large and long-tailed, encompassing millions of entities, many of which are rare or entirely unseen during training. Second, existing visual classifiers treat entities as isolated labels, ignoring the wealth of semantic relationships and factual knowledge that exist between entities. These limitations hinder generalization, especially in zero-shot settings, where the model must recognize entities it has never encountered during training. Third, a key drawback of the two-step generative approaches is the information loss incurred when converting rich visual content into simplified textual labels, leading to semantic ambiguity (Sevgili et al. 2022; Bouarroudj, Boufaïda, and Bellatreche 2022) when conducting BM25 search algorithms based on the predicted textual labels to find the final entity. Entities with similar or identical textual labels may represent fundamentally different concepts. For instance, “Mercury” can denote either the innermost planet of our Solar System or the liquid metal with chemical symbol Hg (atomic number 80), showing how simple text matching cannot distinguish these distinct meanings.
We propose that open-domain visual entity recognition should move beyond superficial recognition toward semantic-level understanding. This involves abstracting both images and textual description into a shared conceptual space with rich structural knowledge among entities (cf. Figure 1b).
(a)(b)Refer to caption Entity ID: Q160659 Name: London Eye Q: What is the wheel here? Entity: Q84 Name: London Q: What is the city in the picture? Entity: Q456960 Name: Westm. Bridge Q: What is the bridge in the front? Entity: Q41225 Name: Big Ben Q: What is clock tower on the left? Entity: Q62408 Name: Palace of Westm. Q: What is palace near the Big Ben? Refer to captionText SpaceImage SpaceSemantic SpaceLondonPalace of Westm.Big BenLondon Eye Ferris Wheel Wheel Tourist Attract. Westm. BridgeRoad. BridgeBridgeP276P276P276P31P31P31P31P31P361P279P279 P276 = Location P31 = Instance of P361 = Part of P279 = Sub Class of Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption The Palace of Westminster is the meeting place of the Parliament of the United King… Big Ben is the nickname for the Great Bell of the Great Clock of Westm… The London Eye, originally the Millennium Wheel, is a cantilevered observation wheel on… Interclass Dist.Intraclass Dist.Modality Gap
Figure 1: (a) The open-domain visual entity recognition task aims to link an image and a text query to a corresponding Wikidata entity. The text query expresses the intent to the image. (b) Both visual and textual inputs can be mapped to a shared semantic space enriched with structured knowledge, such as entity relations (e.g., instance-of, part-of, subclass-of) from Wikidata. Leveraging this additional knowledge source enables contextual reasoning and effective disambiguation of entities.
In this view, recognition is not simply matching visual features to entity names, but of aligning image content with structured, contextualized knowledge about the entities depicted. Available knowledge graphs (KGs), such as Wikidata, which offers a rich and structured representation of hierarchical and association relations for millions of real-world entities, can be leveraged to achieve this.
We present Knowledge-Guided Contrastive Learning (KnowCoL), an approach for open-domain visual entity recognition. Input images and candidate entities’ descriptions and/or lead images are projected into a shared semantic embedding space, structured with the prior knowledge from the KG. This enables zero-shot recognition by allowing the model to generalize from seen to unseen entities based on semantic similarity, while also supporting entity disambiguation via knowledge-informed representations. KnowCoL is based on a dual-encoder paradigm, eliminating the information loss in the process of conversion in the two-step approaches. We evaluate our approach on the OVEN benchmark. Our contributions are as follows:
  1. 1.
    We present a dual-encoder approach for open-domain visual entity recognition task that leverages external knowledge from both Wikidata and Wikipedia to enable a better representation of images and text descriptions.
  2. 2.
    We investigate the impact of different forms of external knowledge (including lead images) and different relations between entities on the recognition performance.
  3. 3.
[…]
Preliminary
Contrastive Loss
Contrastive loss is a widely-used objective in representation learning, particularly in contexts such as metric learning, image-text alignment, and self-supervised learning. Given two sets of embeddings $A=\{a_{i}\}_{i=1}^{N}$ and $B=\{b_{i}\}_{i=1}^{N}$, where $(a_{i},b_{i})$ is a positive (matching) pair and $\forall(a_{i},b_{j}),j\neq i$, is a negative pair, we define the contrastive loss as:
$\ell(A,B)=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\left(\text{sim}(a_{i},b_{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(a_{i},b_{j})/\tau\right)}\text{,}$ (1)
where $\text{sim}(\cdot,\cdot)$ denotes the similarity function, $\tau$ is a temperature parameter controlling the softness of the similarity distribution, and $N$ is the set size. Common similarity functions include cosine similarity and Euclidean distance-based similarity. Specifically, the symmetric contrastive loss can be defined as:
$\ell_{\text{sym}}(A,B)=\frac{1}{2}\left(\ell\left(A,B\right),\ell\left(B,A\right)\right)\text{,}$ (2)
which encourages mutual alignment and it is widely used in various models like CLIP.
Wikidata Knowledge Graph
Wikidata is a collaboratively-maintained, multilingual knowledge graph that can be formalised as the directed graph $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$, where $\mathcal{E}$ is the set of entities (called items and uniquely referenced by “Q-identifiers” (QIDs), e.g., Q42 for Douglas Adams), $\mathcal{R}$ is the set of relation types (called properties and referenced by ”P-identifiers“ (PIDs), e.g., P31 for instance of), and $\mathcal{T}\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ is the set of knowledge triples $(e_{1},r,e_{2})$ stating that subject entity $e_{1}$ is linked to object entity $e_{2}$ via property $r$. As of 2025, Wikidata KG contains more than 100 millions entities, 12,000 properties, and 10 billions of triples and is growing daily through human and bot contributions.
Wikipedia Knowledge Base
Wikipedia is a collaboratively edited, multilingual encyclopedia whose every article is linked to a unique Wikidata QID. This tight coupling lets each page serve as a multimodal hub: rich encyclopedic text is paired with lead images that provide both descriptive and visual context for the corresponding entity in the Wikidata KG.
Related Work
Zero-Shot Classification
Open-world recognition is closely related to zero-shot learning (ZSL), where models must recognize classes with no training images by relying on auxiliary information.
Early approaches (Farhadi et al. 2009; Lampert, Nickisch, and Harmeling 2009) introduced human-defined attributes as intermediate semantic representations for recognition of unseen classes. Later methods moved toward unsupervised semantic embeddings derived from text corpora (Socher et al. 2013; Frome et al. 2013; Norouzi et al. 2014). These dual encoder or projection-based approaches have proven that aligning images with distributed text representations enables zero-shot recognition at scale. Vision-language pretraining has greatly advanced this paradigm. A well-known example is CLIP (Radford et al. 2021) that is trained on 400 million image-text pairs to learn a joint embedding space for images and natural language. By encoding an image and a candidate label (or description) into the same space, CLIP can directly measure their compatibility, essentially performing zero-shot classification by ranking labels. CLIP has wide applications in visual inspection (Jeong et al. 2023; Sadikaj et al. 2025) and robotics (Zhou et al. 2023, 2024a; Yao et al. 2025).
Another line of zero-shot research uses large-scale generative models to bridge the gap to unseen classes. Instead of directly mapping images to text semantics, these generative approaches output the entity names based on given images and then link the entity names to the categories (two-step approach). For instance, Hu et al. (2023) uses PaLI and BLIP-v2 (Li et al. 2023a) on the OVEN benchmark to make the model predict the entity name and then use the BM25 search algorithm to find the closed label in the label space of Wikidata ids. Instead of generating entity names, GER-ALD (Caron et al. 2024) generates the semantic and discriminative “code” to identify entities. Auto-VER (Xiao et al. 2024) combines contrastive learning with generative models to enhance the ability to distinguish similar entities within a vast label space. Unlike dual-encoder approaches, two-step generative methods output text conditioned on images and then match them to labels in the label space. This introduces semantic ambiguity, as generated outputs may be correct lexically but vague or incomplete semantically.
Knowledge Graphs for Visual Entity Recognition
Many studies have explored incorporating knowledge into vision-and-language tasks (Monka, Halilaj, and Rettinger 2022), including visual question answering (Chang et al. 2022; Chen et al. 2021) and entity-aware image captioning (Biten et al. 2019). For the visual entity recognition task, Wang, Ye, and Gupta (2018) and Kampffmeyer et al. (2019) use semantic embeddings and categorical relationships from KGs through graph convolution networks for zero-shot prediction. Li et al. (2023b) augments few-shot image recognition by introducing auxiliary semantic prior knowledge and propagating knowledge among categories via a semantic-visual mapping. KG-NN (Monka et al. 2021) combines prior knowledge encoded in KGs with visual representations to enhance generalization under distribution shifts, and KGV (Zhou et al. 2024b) demonstrates that the integration of richer multi-modal priors can further improve performance. Instead of relying on small domain-specific knowledge graphs, our approach is the first to leverage a large-scale knowledge graph for open-domain visual entity recognition, demonstrating strong scalability potential.
Methodology
FuserRefer to caption LP1 Refer to caption Image Encoder Refer to caption LP2 Refer to caption Text Encoder Refer to captionRefer to caption What’s the wheel in this image? LP1 Refer to caption Image Encoder Refer to caption LP2 Refer to caption Text Encoder Refer to caption The London Eye is a cantilevered observation … Refer to captionRefer to caption Lookup Table Refer to caption London Eye Q160659 London(Q84) Wheel(Q446) Align. Loss Proxy Loss KE Loss Image Emb.Text Emb.Fused Emb.Node Emb.Refer to captionTrainableRefer to captionFrozen Attraction
Figure 2: Framework Overview. Given an input image and text query (left), CLIP encoders extract image and text features. LP1 and LP2 present linear project layers for fine-tuning. A fuser module combines these to represent the intended entity. On the right, prior knowledge of the entity (Q160659 – London Eye)—including sample images, textual descriptions, and knowledge graph structure—is embedded using CLIP’s image and text encoders, as well as knowledge graph embedding (KGE) method. A proxy loss function is employed to align the semantic space (KGEs), visual space (image embeddings), and textual space (text embeddings). Additionally, a knowledge embedding (KE) loss is applied to learn KGEs from the structured knowledge. Finally, an alignment loss is used to align the different input representation with the corresponding entity embedding.
Our goal is to integrate the rich structured knowledge from Wikidata into the training pipeline. To achieve this, we align knowledge graph embeddings (KGEs), image embeddings, and text embeddings within a shared latent space, enabling the structured knowledge captured by the KGEs to be effectively injected into the latent space’s representation. A framework overview can be seen in Figure 2.
Problem Definition
Let $\mathcal{E}$ be the set of Wikidata entities identified by QIDs. Define the input space as $\mathcal{X}=\mathcal{I}\!\times\!\mathcal{L}$, where $\mathcal{I}$ denotes the domain of RGB images and $\mathcal{L}$ is the linguistic (text) space. The Open-domain Visual Entity Recognition (OVEN) task seeks a function $f:\mathcal{X}\rightarrow\mathcal{E}$ which maps each image-text pair ($x^{p}$, $x^{t}$) to the unique entity $e\in\mathcal{E}$. Every entity $e$ is associated with a Wikipedia-derived record $(t_{e},\,I_{e})\in\mathcal{K}\subseteq\mathcal{L}\times 2^{\mathcal{I}}$, where $t_{e}\in\mathcal{L}$ is its encyclopaedic description and $I_{e}\subset\mathcal{I}$ is the finite set of lead images illustrating the corresponding entity. Given a labeled training dataset $\{(x_{i},e_{i})\}_{i=1}^{N}\subset\mathcal{X}\times\mathcal{E}$, the goal is to learn a function $f$ that minimizes the number of misclassifications.
Embeddings
Knowledge Graph Embeddings
To model the knowledge existing in the knowledge graph $\mathcal{G}$, we adopt embedding methods to inject the knowledge into the latent space. We define knowledge graph embeddings as mappings:
$\phi:\mathcal{E}\to\mathcal{Z},\quad\psi:\mathcal{R}\to\mathbb{R}^{d_{r}}\text{,}$
where $\mathcal{Z}\subseteq\mathbb{R}^{d_{e}}$ is the latent space with dimension $d_{e}$. $\phi(e),\psi(r)$ are the node and relation embeddings, respectively. Common knowledge graph embedding methods define score function $f_{s}(e_{1},r,e_{2})$ to measure the plausibility of a triple $(e_{1},r,e_{2})\in\mathcal{T}$.
Image Embeddings
Let $\mathcal{I}$ be the space of input images. We define an image mapping function as:
$f_{\mathbf{\theta}}:\mathcal{I}\to\mathcal{Z}\text{,}$
where $f_{\mathbf{\theta}}$ is represented by a deep neural network with parameters $\mathbf{\theta}$. Here, we leverage the pretrained frozen CLIP image encoder plus a trainable linear project layer to be $f_{\mathbf{\theta}}$.
Text Embeddings
Let $\mathcal{L}$ be the linguistic (text) space of input text descriptions. We define a text mapping function:
$f_{\mathbf{\lambda}}:\mathcal{L}\to\mathcal{Z}\text{,}$
where $f_{\mathbf{\lambda}}$ is represented by a deep neural network with parameters $\mathbf{\lambda}$. We leverage the pretrained frozen CLIP text encoder plus a trainable linear project layer to be $f_{\mathbf{\lambda}}$.
We employ a dual-encoders approach to align the semantic representation of the input image and text query with that of the corresponding entity, which includes its lead image(s) and textual description. We structure our discussion into two parts: (1) encoding the input image and text query, and (2) encoding the multi-modal information of entities from the knowledge graph.
Input Image and Text Query Encoding
To fuse the information of the input image $x^{p}$ and the text query $x^{t}$, we employ an extra encoder
$f_{\mathbf{\gamma}}:(\mathcal{Z},\mathcal{Z})\to\mathcal{Z}$
with parameters $\mathbf{\gamma}$ to to fuse the image embedding $f_{\mathbf{\theta}}(x^{p})$ and text embedding $f_{\mathbf{\lambda}}(x^{t})$. The embedding, combining the information from the input image and text query, can be written as
$\mathbf{z}^{\text{input}}=f_{\mathbf{\gamma}}(f_{\mathbf{\theta}}(x^{p}),f_{\mathbf{\lambda}}(x^{t}))$ (3)
Entity’s Multi-modal Information Encoding
For each entity $e$, $t_{e}$ and $I_{e}$ represent its text description and set of lead image(s), respectively. The entity text embedding can be written as
$\mathbf{z}^{\text{entityText}}=f_{\lambda}(t_{e})\text{,}$ (4)
and the image embedding can be written as
$\mathbf{z}^{\text{entityImage}}=\begin{cases}\frac{1}{|I_{e}|}\sum_{a\in I_{e}}f_{\theta}(a)&\text{if }I_{e}\neq\emptyset\\ \mathbf{z}^{\text{entityText}}&\text{if }I_{e}=\emptyset\end{cases}\text{.}$ (5)
If no lead image is available for a given entity, we assign $\mathbf{z}^{\text{entityText}}$ to $\mathbf{z}^{\text{entityImage}}$. This ensures a smooth fusion of visual and textual information, as discussed in the next section.
Knowledge-guided Contrastive Learning
For given input batch $\{(x^{p}_{i},x^{t}_{i})\}_{i=1}^{N_{b}}$, where $x^{p}_{i}$ and $x^{t}_{i}$ are the $i$-th input images and its corresponding text query, $N_{b}$ is the batch size, the corresponding ground truth label set is denoted as $\{e_{i}\}_{i=1}^{N_{b}}$. We define three losses, namely alignment loss, proxy loss, and knowledge graph loss for the training process. Figure 3 shows the learning visualization.
Alignment Loss
The alignment loss aims to align the joint representation of input images and text queries $\{\mathbf{z}^{\text{input}}_{i}\}_{i=1}^{N_{b}}$ with their corresponding entity representations $\{\phi(e_{i})\}_{i=1}^{N_{b}}$. Specifically, We define the alignment loss using a contrastive objective that pulls matched embeddings closer together while pushing unmatched embeddings among the batch apart. It can be defined as
$\mathcal{L}_{a}=\ell_{\text{sym}}\left(\{\mathbf{z}^{\text{input}}_{i}\}_{i=1}^{N_{b}},\{\phi(e_{i})\}_{i=1}^{N_{b}}\right)\text{,}$ (6)
where $\ell_{\text{sym}}(\cdot,\cdot)$ represents the contrastive loss function define in Equation 2.
Proxy Loss
The proxy loss is defined to align the node embeddings $\{\phi(e_{i})\}_{i=1}^{N_{b}}$ with their corresponding multi-modal representations $\{\mathbf{z}^{\text{entity}}_{i}\}_{i=1}^{N_{b}}$. This objective ensures that the node embeddings capture semantic information from both image and text modalities. It can be defined as:
$\begin{split}\mathcal{L}_{p}=\frac{1}{2}\ell_{\text{sym}}\left(\{\phi(e_{i})\}_{i=1}^{N_{b}},\{\mathbf{z}^{\text{entityText}}_{i}\}_{i=1}^{N_{b}}\right)+\\ \frac{1}{2}\ell_{\text{sym}}\left(\{\phi(e_{i})\}_{i=1}^{N_{b}},\{\mathbf{z}^{\text{entityImage}}_{i}\}_{i=1}^{N_{b}}\right)\end{split}$ (7)
We refer to this as the proxy loss because the node embeddings serve as proxies for capturing the multi-modal semantic information of the entities. Figure 3 shows that the proxy loss gathers image and text embeddings with their corresponding node embeddings.
Encoder Refer to caption “London Eye…” Refer to caption “Wheel is a…” Step 0 Q160659Q446 P31 Step N Q160659Q446P31Image Emb.Text Emb.Node Emb. Rel. Emb.
Figure 3: Node embeddings act as proxies, aligned with their corresponding image and text embeddings. Knowledge (from the KG) is captured through relational embeddings, connecting node embeddings in the latent space. Q160659, Q446, P31 represent London Eye, Wheel, and Instance of.
Knowledge Graph Embedding Loss (KE Loss)
The knowledge graph loss aims to capture the structural knowledge associated with entity $e_{i}$. We first extract the triplet set associated with the entity $e_{i}$ by $\mathcal{T}_{e_{i}}=\{(h,r,t)\in\mathcal{T}\mid h=e_{i}\lor t=e_{i}\}$. For each triplet $(h,r,t)\in\mathcal{T}_{e_{i}}$, we define $\mathcal{T}_{(h,r,t)}$ as the set of negative samples generated by corrupting either the head or tail entity in the positive triplet:
$\mathcal{T}_{\text{(h,r,t)}}=\{(h^{\prime},r,t)\mid h^{\prime}\in\mathcal{E}\setminus h\}\cup\{(h,r,t^{\prime})\mid t^{\prime}\in\mathcal{E}\setminus h\}\text{.}$ (8)
Based on that, the knowledge graph loss can be defined as
$\mathcal{L}_{\text{KE}}$ $=\sum_{i=1}^{N_{b}}\frac{1}{|\mathcal{T}_{e_{i}}|}\sum_{(h,r,t)\in\mathcal{T}_{e_{i}}}\frac{1}{|\mathcal{T}_{(h,r,t)}|}$ (9)
$\quad\left[-\log\frac{\exp(f_{s}(h,r,t)/\tau)}{\sum_{(h^{\prime},r,t^{\prime})\in\mathcal{T}_{(h,r,t)}}\exp(f_{s}(h^{\prime},r,t^{\prime})/\tau)}\right]\text{,}$
where $\tau$ is a temperature and $f_{s}(\cdot)$ is the score function defined by
$f_{s}(h,r,t)=\frac{(\phi(h)+\psi(r))^{\top}\phi(t)}{\|\phi(h)+\psi(r)\|_{2}\,\|\phi(t)\|_{2}}\text{,}$ (10)
which is the cosine similarity between $\phi(h)+\psi(r)$ and $\phi(t)$. Here, we follow a translation-based KGE approach with a cosine similarity distance metric. The final loss for optimization can be defined as:
$\mathcal{L}=\mathcal{L}_{a}+\beta_{1}\mathcal{L}_{p}+\beta_{2}\mathcal{L}_{\text{KE}}\text{,}$ (11)
where $\beta_{1}$ and $\beta_{2}$ are two hyperparameters for balancing.
Inference
Given an input image $x^{p}$ and a text query $x^{t}$, we first obtain the input embedding $\mathbf{z}^{\text{input}}$ as defined in Equation 3. The corresponding entity answer $e^{*}$ is then identified through the following inference step:
$e^{*}=\arg\max_{e_{j}\in\mathcal{E}}\text{sim}\left(\mathbf{z}^{\text{input}},\,\frac{1}{2}\left(\mathbf{z}^{\text{entityText}}_{j}+\mathbf{z}^{\text{entityImage}}_{j}\right)\right)\text{,}$ (12)
where $\mathbf{z}^{\text{entityText}}_{j}$ and $\mathbf{z}^{\text{entityImage}}_{j}$ denote the textual and visual embeddings of entity $e_{j}$, respectively, as defined in Equations 4 and 5.
Experiments
In our experiments, we investigate whether incorporating knowledge from large-scale KGs and KBs can enhance open-domain visual entity recognition. We analyze the impact of different types of prior knowledge, fusion methods, and KGE techniques. In addition, ablation studies on various hyperparameter values are also provided.
Experiment Settings
Paradigm Model Venues Parameters (B) Pre-train Dataset Seen Unseen HM
Dual-Encoder
CLIP (Radford et al. 2021) ICML2021 0.4 OpenAI 5.6 4.9 5.2
CLIPFusion (Hu et al. 2023) ICCV2023 0.9 OpenAI 33.6 4.8 8.4
CLIP2CLIP (Hu et al. 2023) ICCV2023 0.9 OpenAI 12.6 10.5 11.4
Two-Step
Generative
BLIP-v2 (Li et al. 2023a) ICML2023 12.2 - 8.6 3.4 4.9
PaLI-3B (Hu et al. 2023) ICCV2023 3.0 WebLI 19.1 6.0 9.1
PaLI-17B (Hu et al. 2023) ICCV2023 17.0 WebLI 28.3 11.2 16.0
Git-Large (Caron et al. 2024) CVPR2024 0.4 WebLI 17.6 4.3 4.0
GER-ALD (Caron et al. 2024) CVPR2024 0.4 LAION 29.1 17.7 22.7
Auto-Ver-7B (Xiao et al. 2024) ECCV2024 7.0 - 62.8 16.0 25.5
Auto-Ver-14B (Xiao et al. 2024) ECCV2024 14.0 - 65.0 18.6 28.9
Knowledge-Guided
Dual-Encoder
KnowCoL-bigG(ours) - 2.0 LAION 41.8 36.1 38.8
Table 1: Comparison with current state-of-the-art approaches on OVEN entity test split. We evaluate the harmonic mean (HM) of the seen and unseen splits (top1 accuracy) after fine tuning on OVEN training set. It serves as our main metric. The numbers of baselines are taken from papers (Caron et al. 2024; Hu et al. 2023; Xiao et al. 2024). The total parameters and pre-training datasets of the models are given for comparison.
OVEN Dataset
The OVEN dataset contains 6,063,945 training samples, which combines 14 image recognition datasets. Models are evaluated on the test splits, and performance is reported using the harmonic mean (HM) of top-1 accuracy for two types of entities: seen entities, which appear in the OVEN training set, and unseen entities, which are not present during training. Specifically, the test split includes 15,888 entities (8,355 seen and 7,533 unseen). Our evaluation is conducted solely on the OVEN benchmark, comprising 14 diverse image recognition datasets, thus indicating the generalizability of our approach. The evaluation metric is top-1 accuracy for seen and unseen entities, and the harmonic mean (HM) of them.
Training details
We extract a subgraph from Wikidata containing 32,122 entities and 501 relation types. For training, we use the AdamW optimizer with a learning rate of 0.001, batch size of 4096, and weight decay of 0.0001. Our approach includes three model variants based on different OpenCLIP image encoders: ViT-L/14, ViT-H/14, and ViT-bigG/14. The temperature $\tau$ for contrastive learning is fixed at 0.07 during fine-tuning. The hyperparameters $\beta_{1}$ and $\beta_{2}$ are both set to 1. We fuse information from the input image and text query (Equation (3)), using a simple addition operation. More details are provided in the Appendix.
Baselines
Dual Encoders
Dual encoders approaches consist of two encoders, each encoding a dedicated modality (e.g., images and text) into the same latent space. The visual recognition is conducted by searching the nearest neighbors of prototypes in the latent space for a given input image. Representative models include CLIP, CLIPfusion, and CLIP2CLIP.
Auto-regressive Captioning
Another approach uses auto-regressive models to generate captions for the given images. Representative vision-language models (VLMs) such as PaLI, GiT-Large, GER-ALD, Auto-VER follow this paradigm. These models generate a descriptive caption or name for the visual entity and then match it to the closest label in the predefined label space for final prediction.
Comparison with the state of the art
[…]
Analysis and Ablation Studies
Here, we analyse the impact of model size, prior knowledge types, fusion strategy of the text query and input image, various KGE methods, and hyperparameters - $\beta_{1}$, $\beta_{2}$, latent space dimension $d_{e}$, and temperature $\tau$. The detailed setting can be found in the Appendix section experiment setting.
Impact of Model Size
Model Backbone Para.(B) Seen Unseen HM
CLIP ViT-L-14 0.4 5.6 4.9 5.2
CLIPFusion ViT-L-14 0.9 33.6 4.8 8.4
CLIP2CLIP ViT-L-14 0.9 12.6 10.5 11.4
KnowCoL-L ViT-L-14 0.4 34.3 29.1 31.5
KnowCoL-H ViT-H-14 0.9 38.5 33.4 35.8
KnowCoL-bigG ViT-bigG-14 2.0 41.8 36.1 38.8
Table 2: Comparison of model sizes. ViT-H-14 and ViT-bigG-14 are used for our KnowCoL-H and KnowCoL-bigG models. ViT-L-14 is used for CLIP, CLIPFusion, CLIP2CLIP, and KnowCoL-L. Note that CLIPFusion and CLIP2CLIP utilize two ViT-L-14 backbones.
As Table 2 demonstrates, we evaluate our model with different CLIP backbones. The backbones ViT-L-14, ViT-H-14, and ViT-bigG-14 are used for KnowCoL-L, KnowCoL-H, and KnowCoL-bigG, respectively. The performance across both seen and unseen entities consistently improves as model size increases.
We highlight that KnowCoL-L, CLIP, CLIPFusion, and CLIP2CLIP use the same backbone ViT-L-14, but KnowCoL-L significantly outperforms all. While CLIPFusion and CLIP2CLIP achieve harmonic means of 8.4% and 11.4% respectively, KnowCoL-L reaches 31.5%, showing over 3 times improvement. This demonstrates that external knowledge significantly improves zero-shot generalization without increasing model size.
0.10.10.50.511303035354040HMKnowCoL-LKnowCoL-H
(a) Hyperparameter $\beta_{1}$
0.10.10.50.511303035354040HMKnowCoL-LKnowCoL-H
(b) Hyperparameter $\beta_{2}$
51251276876810241024303035354040HMKnowCoL-LKnowCoL-H
(c) Latent Space Dim $d_{e}$
0.050.050.070.070.10.1303035354040HMKnowCoL-LKnowCoL-H
(d) Temperature $\tau$
Figure 4: Ablation Studies for hyperparameter $\beta_{1}$, $\beta_{2}$, latent space dimension $d_{e}$, and temperature $\tau$. HM indicates the harmonic mean of accuracies of seen and unseen entities. - represents the default setting of KnowCoL approach.
Comparison of Knowledge Types
Model Seen Unseen HM
[…]
w/o Lead images 32.5 27.9 30.3
w/o KG Hierarchical Knowledge 32.1 25.3 28.3
Table 3: Impact of different types of knowledge. We investigate the impact of structural knowledge, namely the sample images in the KB and the hierarchical relations in the KG.
Table 3 demonstrates the importance of different knowledge types in the KnowCoL-L model. The complete KnowCoL-L model achieves the best score of 31.5% HM. Removing lead images (w/o Lead images) reduces performance to 32.5% (seen), 27.9% (unseen), and 30.3% (HM), highlighting the role of visual context. Omitting hierarchical KG knowledge of ‘Instance of’, ‘subclass of’, and ‘parent taxon’ further lowers scores to 32.1% (seen), 25.3% (unseen), and 28.3% (HM), significantly impairing unseen entity recognition.
Comparison of Fusion Functions
Fusion Method Layer Number Seen Unseen HM
[…]
Concatenation + MLP 1 37.5 26.8 31.2
Concatenation + MLP 2 45.4 21.0 28.7
TE 2 54.4 16.6 25.5
TE + IP 2 56.7 17.9 27.2
Table 4: Comparison of fusion functions for combining image and text embeddings of KnowCoL-L. Our default fusion method is addition. ‘TE’ refers to the Transformer Encoder used in (Hu et al. 2023). ‘TE + IP’ is inspired by LLaVA (Liu et al. 2023). In this variant, the fusion method also incorporates local image patch tokens from the ViT encoder.
As Table 4 shows, the default Addition method provides balanced performance, achieving the best harmonic mean 31.5%. By using the multi-layer perceptron (MLP) method, the performance for seen entities improves up to 37.5%, while for unseen entities decreases to 26.8%. Using transformer encoder (TE) methods with or without local image patch embeddings, significantly boost the accuracy on seen entities (54.4% and 56.7%), but degrade unseen entity recognition (16.6% and 17.9%). The trend shows that more complex fusion methods, containing more layers or a more complex structure, lead to overfitting on seen entities and sacrificing zero-shot generalization to unseen entities. Additionally, by introducing local image patch tokens (TE + IP), performance on both seen (56.7%) and unseen entities (17.9%) improves compared to the TE-only method. This improvement likely occurs because local image patch tokens provide fine-grained visual details and enhance recognition ability.
Comparison of KGE Methods
KGE Methods Distance Measure Seen Unseen HM
[…]
TransE Euclidean Dis. 33.7 28.5 31.0
TransH Cosine Similarity 34.4 28.6 31.2
DistMult - 33.8 28.5 30.9
Table 5: Comparison of KGE methods for incorporating structural knowledge. Our default approach uses TransE combined with a cosine similarity metric. We compare this against alternative configurations, including TransE with Euclidean distance, TransH, and DistMult.
Table 5 compares various knowledge graph embedding methods and distance measures. Our default choice TransE (Bordes et al. 2013) with cosine similarity, achieves the best overall performance. This setup aligns well with CLIP, which also uses cosine similarity in its contrastive learning. When replacing the cosine similarity with Euclidean distance, the HM drops from 31.5% to 31.0%. TransH (Wang et al. 2014) introduces more representational flexibility by allowing entity embeddings to vary across different relations. This additional modeling capacity slightly improved seen performance to 34.4% but comes at the cost of lower unseen accuracy to 28.6%. The DistMult method (Yang et al. 2015) shows the lowest performance, suggesting it is less effective for integrating structural knowledge from the KG into the visual-textual space.
Hyperparameters
As Figure 4 shows, we conduct experiments varying the hyperparameters $\beta_{1}$, $\beta_{2}$, latent space dimension $d_{e}$, and contrastive learning temperature $\tau$. Figure 4a shows that increasing the weight of the proxy loss $\beta_{1}$, aligning node embeddings with multi-modal prior knowledge, enhances model performance. Figure 4b demonstrates that increasing $\beta_{2}$, the weight of KE loss incorporating structured prior knowledge from Wikidata, also improves performance. Additionally, we identify that the larger model (KnowCoL-H) with more prior knowledge can benefit more from the inclusion of structured prior knowledge. Figure 4c illustrates that larger models require higher dimensional latent spaces to represent richer features. The temperature 0.07 performs best in our task, as shown in Figure 4d.
Conclusion and Future Works
We propose KnowCoL, a novel approach to improve open-domain visual entity recognition by incorporating multi-modal prior knowledge, namely structured knowledge, textual and visual priors. KnowCoL significantly enhances visual entity recognition performance, especially for zero-shot recognition of unseen entities. Future work includes exploring non-Euclidean spaces like hyperbolic and hyperspherical spaces to better capture the hierarchical relationships and semantic structures inherent in knowledge graphs. Investigating these non-Euclidean geometries could further improve representation learning, potentially leading to more accurate and efficient entity recognition and generalization.
Acknowledgement
This work was partially funded by the European Union’s Horizon RIA research and innovation programme under grant agreement No. 101092908 (SMARTEDGE). The authors also thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Hongkuan , Yuqicheng Zhu, and Jingcheng Wu.
Appendix
The following sections present detailed descriptions of the datasets and training procedures for each experimental scenario. We outline the experimental settings for each variant of our approach, including different fusion strategies for image and text embeddings, the score functions employed by various knowledge graph embedding (KGE) methods, and key hyperparameters. Additionally, we introduce our approach for extracting efficient subgraphs from Wikidata. Finally, we provide representative examples of multi-modal prior knowledge utilized in our KnowCoL approach for a clearer overview. The pseudo code for KnowCoL algorithms are listed in Algorithm 2 and  3.
Experiment Setting
Dataset
OVEN dataset consists of 14 existing datasets, grounding their labels to Wikidata ids. These 14 datasets are originally created for image recognition/retrieval, and visual question answering. They are ImageNet21k-P (Ridnik et al. 2021), iNaturalist2017 (Horn et al. 2018), Cars196 (Krause et al. 2013), SUN397 (Xiao et al. 2010), Food101 (Bossard, Guillaumin, and Gool 2014), Sports100, Aircraft (Maji et al. 2013), Oxford Flower (Nilsback and Zisserman 2008), Google Landamarks v2 (Weyand et al. 2020), VQA v2 (Goyal et al. 2019), Visual7W (Zhu et al. 2016), Visual Genome (Krishna et al. 2017), OK-VQA (Marino et al. 2019), and Text-VQA (Singh et al. 2019).
Image and Text Query Fuser
In this section, we elaborate the different fusion function in detail, namely addition, multi-layer perception (MLP), transformer encoder (TE), and transformer encoder with image patch embeddings (TE+IP).
Addition
The default fuser we used is simple addition operation, as it shows the best generalization ability, especially for unseen entities. The fused embedding is defined as follows:
$\mathbf{z}^{\text{input}}=f_{\mathbf{\theta}}(x^{p})+f_{\mathbf{\lambda}}(x^{t})\text{,}$ (13)
where $f_{\theta}(\cdot)$ and $f_{\lambda}(\cdot)$ are the image encoder and text encoder, respectively.
Multi-layer Perception (MLP)
The variant Multi-Layer Perceptron (MLP) fuser first concatenates the input image embedding and the text query embedding, each of dimension $d_{e}$, into a single embedding of dimension $2\times d_{e}$. An MLP encoder then reduces this embedding dimension from $2\times d_{e}$ back down to $d_{e}$. When the fusion layer is set to 1, the MLP consists of a single linear layer (input dimension $2\times d_{e}$, output dimension $d_{e}$) followed by a ReLU activation function. When the fusion layer is set to 2, an additional intermediate layer of dimension $d_{e}$ is introduced, resulting in two linear layers, each followed by a ReLU activation. The fused embedding is represented as follows:
$\mathbf{z}^{\text{input}}=\text{MLP}\left(\text{Concat}\left(f_{\theta}(x^{p}),f_{\lambda}(x^{t})\right)\right)\text{,}$ (14)
where $\text{MLP}(\cdot)$ is the function for multi-layer perception and $\text{Concat}(\cdot)$ is the function for concatenation.
Transformer Encoder (TE)
The variant transformer encoder (TE) fuses the image embedding and the text embedding with transformer encoder on top. Both embeddings are fed into the encoder, allowing them to attend to each other through self-attention. The first output embedding from the encoder is used as a global representation, capturing the fused multi-modal information.
Transformer Encoder with Image Patch Embeddings (TE + IP)
Compared to the previous approach, the local image patch embeddings from the image encoder are also fed into the transformer encoder. Similarily, the first output embedding from the encoder is used as a global representation, capturing the fused information.
KGE Method Settings
Our default setting adopts the TransE method, replacing the original Euclidean distance metric with cosine similarity. The motivation for this change is that CLIP is pre-trained using cosine similarity, and adopting the same metric ensures better alignment with the embedding space learned by CLIP. Experimental results also show that using cosine similarity improves performance. Instead of using the original margin-based formulation for negative samples, we adopt a cross-entropy-based approach, which is similar to that used in CLIP, in order to ensure that the KE loss is on the same scale as the main alignment loss. The details of other settings are given as follows.
TransE + Euclidean
In this setting, we adopt the original TransE method, using Euclidean distance and margin-based calculation for negative samples. The KE loss is defined as follows:
$\mathcal{L}_{\text{KE}}$ $=\sum_{i=1}^{N_{b}}\frac{1}{|\mathcal{T}_{e_{i}}|}\sum_{(h,r,t)\in\mathcal{T}_{e_{i}}}\frac{1}{|\mathcal{T}_{(h,r,t)}|}\sum_{(h^{\prime},r,t^{\prime})\in\mathcal{T}_{(h,r,t)}}$ (15)
$\quad\times\!\max\!\Bigl(0,\,\epsilon+f_{s}({h},{r},{t})-f_{s}({h}^{\prime},{r},{t}^{\prime})\Bigr)\text{,}$
where $\epsilon$ represents the margin and $f_{s}(\cdot)$ is the score function defined by
$f_{s}(h,r,t)=\|\phi(h)+\psi(r)-\phi(t)\|_{2}\text{.}$ (16)
Here, $\phi(\cdot)$ and $\psi(\cdot)$ are functions which maps nodes and relation types to corresponding embeddings.
TransH + Cosine Similarity
TransH is an extension for TransE, which defines relation-specific hyperplane instead of directly translation in the entity space. It can be realized by altering the score function with
$\begin{split}f_{s}(h,r,t)=\|(\phi(h)-\mathbf{w}_{r}^{\top}\phi(h)\mathbf{w}_{r})+\psi(r)\\ -(\phi(t)-\mathbf{w}_{r}^{\top}\phi(t)\mathbf{w}_{r})\|_{2}\text{,}\end{split}$ (17)
where $\mathbf{w}_{r}$ defines the normal vector of the hyperplane. We still use the cosine similarity and cross entropy-based approaches to conduct contrastive learning.
DistMult
The score function of DistMult is defined by
$\text{score}(h,r,t)\;=\;\phi(h)^{\top}\,\mathrm{diag}(\psi(r))\,\phi(t),$ (18)
where $\mathrm{diag}(\mathbf{r})\in\mathbb{R}^{d\times d}$ is the diagonal matrix with $\psi(r)$ on its diagonal.Similarly, we use the cosine similarity and cross entropy-based approaches to conduct contrastive learning.
Hyperparameters and Training Settings
The hyperparameters used for all three variants of KnowCoL approach, namely KnowCoL-L, KnowCoL-H, KnowCoL-bigG, are listed as follows:
KnowCoL-L
We adopt the OpenCLIP ViT-L/14 image encoder as our backbone and train for 15 epochs with a batch size of 4096. Optimization is performed using AdamW (learning rate $\beta_{1}=1.0$, $\beta_{1}=1.0$, $\beta_{2}=1.0$, weight decay $1\times 10^{-4}$) together with a cosine-annealing learning-rate schedule. We set the embedding dimension $d_{e}$ to 768 and the temperature of contrastive learning $\tau$ to 0.07. In our knowledge-graph module, each entity is associated with up to 50 triplets, and we randomly sample 25 negative triplets per entity. The model was trained using 4 NVIDIA H100 GPUs, each with 80GB of memory, over a total duration of approximately 22 hours. The training was performed using mixed precision (FP16).
KnowCoL-H
We adopt the OpenCLIP ViT-H/14 image encoder as our backbone. The model is trained for 20 epochs with a batch size of $4096$. The AdamW optimizer is employed with a learning rate of $1\times 10^{-3}$, cosine annealing scheduling, and weight decay of $1\times 10^{-4}$. The hyperparameters $\beta_{1}$, $\beta_{2}$, latent dimension $d_{e}$, and contrastive learning temperature $\tau$ are set as $1.0$, $1.0$, $1024$, and $0.07$, respectively. For the knowledge graph, each entity is associated with at most $50$ triplets, and we draw $25$ negative samples per entity. The model is trained with 8 NVIDIA H100 GPUs with 80GB memory. The total training time is around 25 hours. The training was performed using mixed precision (FP16).
KnowCoL-bigG
We adopt the OpenCLIP ViT-bigG/14 image encoder as our backbone. The model is trained for 25 epochs with a batch size of $4096$. The AdamW optimizer is employed with a learning rate of $1\times 10^{-3}$, cosine annealing scheduling, and weight decay of $1\times 10^{-4}$. The hyperparameters $\beta_{1}$, $\beta_{2}$, latent dimension $d_{e}$, and contrastive learning temperature $\tau$ are set as $1.0$, $1.0$, $1024$, and $0.07$, respectively. For the knowledge graph, each entity is associated with at most $50$ triplets, and we draw $25$ negative samples per entity. The model is trained with 8 Nvidia H200 GPUs with 140GB of memory. The total training time is about 30 hours. The training was performed using mixed precision (FP16).
Structure Knowledge Extraction
Wikidata comprises billions of entities and thousands of properties, making it computationally infeasible to utilize the entire knowledge graph for model training. Moreover, the noise within Wikidata poses an additional hurdle, potentially obstructing the learning process and affecting model performance. To mitigate this challenge, we extract a subgraph from Wikidata that encapsulates the essential information required for the OVEN benchmark. We mainly focus on the entity-to-entity relationships among all entities occurring in the OVEN benchmark, plus the hierarchical relations ‘subclass of’ (P279), ‘instance of’ (P31), and ‘parent taxon’(P171) to preserve the structural integrity of the knowledge representation.
Algorithm 1 Subgraph Extraction from Wikidata
1: Given:
  • Original Wikidata KG $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$
  • Dataset entity set $\mathcal{E}_{\text{data}}$
  • Hierarchical relations $p_{1}$ (‘subclass of’), $p_{2}$ (‘instance of’), and optionally $p_{3}$ (‘parent taxon’).
2: Initialize $\mathcal{E}_{\text{data}}^{+}\leftarrow\mathcal{E}_{\text{data}}$
3:for each entity $e\in\mathcal{E}_{\text{data}}$ do
4:  Find superclasses: $S_{e}=\{e^{\prime}\mid(e,p_{1},e^{\prime})\in\mathcal{T}\lor(e,p_{2},e^{\prime})\in\mathcal{T}\}$
5:  Update $\mathcal{E}_{\text{data}}^{+}\leftarrow\mathcal{E}_{\text{data}}^{+}\cup S_{e}$
6:end for
7: Initialize $\mathcal{T}_{\text{sub}}\leftarrow\emptyset$, $\mathcal{R}_{\text{sub}}\leftarrow\emptyset$
8:for each triple $(e_{1},r,e_{2})\in\mathcal{T}$ do
9:  if $e_{1}\in\mathcal{E}_{\text{data}}^{+}$ and $e_{2}\in\mathcal{E}_{\text{data}}^{+}$ then
10:   $\mathcal{T}_{\text{sub}}\leftarrow\mathcal{T}_{\text{sub}}\cup\{(e_{1},r,e_{2})\}$
11:   $\mathcal{R}_{\text{sub}}\leftarrow\mathcal{R}_{\text{sub}}\cup\{r\}$
12:  end if
13:end for
14: Define induced subgraph: $\mathcal{G}_{\text{sub}}=(\mathcal{E}_{\text{data}}^{+},\mathcal{R}_{\text{sub}},\mathcal{T}_{\text{sub}})$
15:return $\mathcal{G}_{\text{sub}}$
The Wikidata knowledge graph is denoted as $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$, where $\mathcal{E}$, $\mathcal{R}$ represents the entity set and predicates set, respectively. $\mathcal{T}\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ is the set of triples ($e_{1}$, $r$, $e_{2}$) where $e_{1}$, $e_{2}\in\mathcal{E}$ and $r\in\mathcal{R}$.
The entity set of the dataset is denoted as $\mathcal{E}$. We first expand it by incorporating all superclasses of the entities in E{E}, as defined by the ‘subclass of’ $p_{1}$ and the ‘instance of’ $p_{2}$ relation in Wikidata. The resulting expanded set, denoted as $E^{+}$, is defined as:
$\mathcal{E}_{\text{data}}^{+}={\mathcal{E}_{\text{data}}}\cup\{e^{\prime}\mid e\in\mathcal{E}_{\text{data }},(e,p_{1},e^{\prime})\in\mathcal{G}\lor(e,p_{2},e^{\prime})\in\mathcal{G}\}\text{,}$ (19)
where $e^{\prime}$ denotes a superclass of entity $e$.
We define the subgraph $\mathcal{G}_{\text{sub}}$ as the induced subgraph containing all entities in $\mathcal{E}^{+}_{\text{data}}$ and all relations that exist between these entities. Formally, the subgraph $\mathcal{G}_{\text{sub}}$ is defined as:
$\mathcal{G}_{\text{sub}}=(\mathcal{E}^{+}_{\text{data}}\mathcal{R}_{\text{sub}},\mathcal{T}_{\text{sub}})$ (20)
where
$\mathcal{R}_{\text{sub}}$ $=\{r\mid\exists e_{1},e_{2}\in\mathcal{E}^{+}_{\text{data}},(e_{1},r,e_{2})\in\mathcal{T}\},$ (21)
$\mathcal{T}_{\text{sub}}$ $=\{(e_{1},r,e_{2})\in\mathcal{T}\mid e_{1},e_{2}\in\mathcal{E}^{+}_{\text{data}}\land r\in\mathcal{R}\}.$
This subgraph $\mathcal{G}_{\text{sub}}$ contains the relations between entities existing in the OVEN benchmark and the hierarchical relations between the entities and their super classes. The pseudo code can be seen in Algorithm 1.
Algorithm 2 Knowledge-Guided Contrastive Learning - Training
1: Given:
  • Dataset $D=\{(x^{p}_{i},x^{t}_{i}),e_{i}\}_{i=1}^{N}$
  • Knowledge Base $\mathcal{K}=\{(t_{e_{i}},I_{e_{i}})\}_{i=1}^{|\mathcal{E}|}$, where $t_{e_{i}}$ is the text description and $I_{e_{i}}$ is the lead images set of entity $e_{i}$
  • Knowledge Graph $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$
  • The model $\mathcal{F}=\{f_{\theta},f_{\lambda},f_{\gamma}\,\phi,\psi\}$ with pre-trained image and text encoders $f_{\theta}$, $f_{\lambda}$, fusion function $f_{\gamma}$, and entity and relation mapping function $\phi,\psi$
2: Initialize projection layers of encoders $f_{\theta}$, $f_{\lambda}$, and fusion $f_{\gamma}$, the other parameters of the encoders remain fixed.
3: Randomly initialize entity embeddings lookup table $\{\phi(e_{i})\}_{i=1}^{|\mathcal{E}|}$ and relation embeddings lookup table $\{\psi(r_{i})\}_{i=1}^{|\mathcal{R}|}$
4:while not converged do
5:  Sample batch $\{(x^{p}_{i},x^{t}_{i},e_{i})\}_{i=1}^{N_{b}}$ from $D$
6:  $\mathbf{z}_{i}^{\text{input}}=f_{\gamma}(f_{\theta}(x_{i}^{p}),f_{\lambda}(x_{i}^{t}))$
7:  $\mathbf{z}_{i}^{\text{entityText}}=f_{\lambda}(t_{e_{i}})$
8:  $\mathbf{z}_{i}^{\text{entityImage}}=\frac{1}{|I_{e}|}\sum_{a\in I_{e}}f_{\theta}(a)$ if $I_{e_{i}}\neq\emptyset$ else $z_{i}^{\text{entityText}}$
9:  Compute alignment loss $L_{a}$, proxy loss $L_{p}$, and knowledge graph embedding loss $L_{KE}$
10:  Compute total loss $L=L_{a}+\beta_{1}L_{p}+\beta_{2}L_{KE}$
11:  Update parameters $\theta,\lambda,\gamma$ based on $L$
12:end while
Algorithm 3 Knowledge-Guided Contrastive Learning - Inference
1: Given
  • Image $x^{p}$
  • Text query $x^{t}$
2: Compute $\mathbf{z}^{\text{input}}=f_{\gamma}(f_{\theta}(x^{p}),f_{\lambda}(x^{t}))$
3: Return predicted entity $e^{*}=\arg\max_{e_{j}\in\mathcal{E}}\text{sim}\left(\mathbf{z}^{\text{input}},\,\frac{1}{2}\left(\mathbf{z}^{\text{entityText}}_{j}+\mathbf{z}^{\text{entityImage}}_{j}\right)\right)$
Additional Experiments
[…]
Methods
Duplicate
Numbers
Query Split
Seen Unseen HM
KnowCoL-L 1 8.5 6.1 7.1
KnowCoL-L 10 11.7 7.3 8.9
KnowCoL-L 50 16.2 8.0 10.7
KnowCoL-H 50 18.5 10.5 13.4
KnowCoL-bigG 50 21.7 12.1 15.6
Table 6: Query Split Results
Multi-modal Prior Knowledge of
Sample Entities
Table 7 presents representative examples of prior knowledge used in the proposed model. Specifically, we provide three illustrative examples each for animals, cars, landmarks, and people. For each example, we include a lead image, a concise textual description, and associated structural knowledge from Wikidata. The structured relationships relevant to these entities, such as instance of’ (P31), subclass of’ (P279), and ‘parent taxon’ (P171), among other significant properties, are detailed in the table. This structured knowledge helps enrich entity representation, facilitating semantic disambiguation and enhancing the generalization capabilities of our KnowCoL framework.
Table 7: Prior Knowledge of Sample Entities: for each entity we show its lead image, a brief textual description, and its structural knowledge as key Wikidata triples (with that entity as head).
Entity Predicate Entity
Name (QID) Lead Image Text Description Name (PID) Name (QID)
Mountain Hare
[Uncaptioned image] The mountain hare (Lepus timidus), also known as blue hare, tundra hare, variable hare, white hare, snow hare, alpine hare, and Irish hare, is a species of Palearctic hare that is largely adapted to polar and mountainous habitats… Instance of (P31) Taxon (Q16521)
Taxon Rank (P105) Species (Q7432)
Parent Taxon (P171) Lepus (Q3830767)
Diel Cycle (P9566) Nocturnal (Q101029366)
Habitat (P2974) Grassland (Q1006733)
Platypus
[Uncaptioned image] The platypus (Ornithorhynchus anatinus), sometimes referred to as the duck-billed platypus, is a semiaquatic, egg-laying mammal endemic to eastern Australia, including Tasmania. The platypus is the sole living representative or … Instance of (P31) Taxon (Q16521)
Taxon Rank (P105) Species (Q7432)
Parent Taxon (P171) Ornithorhyn. (Q3745848)
Endemic To (P183) Australia (Q408)
Characteristic (P1552) Oviparity (Q212306)
Morpho
Menelaus
[Uncaptioned image] The Menelaus blue morpho (Morpho menelaus) is one of thirty species of butterfly in the subfamily Morphinae. Its wingspan is approximately 12 cm (4.7”), and its dorsal forewings and hindwings are a bright, iridescent blue edged with black … Instance of (P31) Taxon (Q16521)
Taxon Rank (P105) Species (Q7432)
Parent Taxon (P171) Morpho (Q645804)
Color (P462) Blue (Q1088)
Has Part(s) (P527) Insect Wing (Q276572)
BMW X5
[Uncaptioned image] The BMW X5 is a mid-size luxury crossover SUV produced by BMW. The X5 made its debut in 1999 as the E53 model. It was BMW’s first SUV. At launch, it featured all-wheel drive and was available with either a manual or … Instance of (P31) Automobile M. (Q3231690)
Manufacturer (P176) BMW (Q26678)
Subclass of (P279) Executive Car (Q1357619)
Subclass of (P279) Car (Q1420)
Subclass of (P279) Road Vehicle (Q1515493)
Tesla Model S
[Uncaptioned image] The Tesla Model S is a battery-electric, four-door full-size car produced by the American automaker Tesla since 2012. The automaker’s second vehicle and longest-produced model, the Model S has both received mixed reviews from critics … Instance of (P31) Automobile M. (Q3231690)
Manufacturer (P176) Tesla, Inc. (Q478214)
Follows (P155) Tesla Roadster (Q210893)
Followed by (P156) Model X (Q1634161)
Subclass of (P279) Electric Car (Q193692)
Volkswagen
Beetle
[Uncaptioned image] The Volkswagen Beetle, officially the Volkswagen Type 1, is a small family car produced by the German company Volkswagen from 1938 to 2003. One of the most iconic cars in automotive history, the Beetle is noted for its distinctive shape… Instance of (P31) Automobile M. (Q3231690)
Instance of (P31) Vehicle M. (Q29048322)
Brand (P1716) Volkswagen(Q246)
Origin Country (P495) Germany (Q183)
Powered by (P516) Gas. Engine (Q502048)
Taj
Mahal
[Uncaptioned image] The Taj Mahal is an ivory-white marble mausoleum on the right bank of the river Yamuna in Agra, Uttar Pradesh, India. It was commissioned in 1631 by the fifth Mughal emperor, Shah Jahan to house the tomb of his beloved wife … Instance of (P31) Tomb (Q381885)
Instance of (P31) Tourist Attr.(Q570116)
Religion (P140) Islam(Q432)
Country (P17) India (Q668)
Architect (P84) Ahmad Lahori (Q2551253)
Potala
Palace
[Uncaptioned image] Potala Palace is the name of a museum in the Tibet Autonomous Region of China previously a palace of the Bö sovereign in Ü, the Dalai Lama, in the dzong-style, in Lhasa, capital of Bod and historically Ü. It was the winter palace of the … Instance of (P31) Palace (Q16560)
Instance of (P31) Tourist Attr.(Q570116)
Has Use (Q71229) Official Resi. (Q481289)
Different From (P1889) State Housing (Q7603702)
Country (P17) China (Q148)
Faroe
Islands
[Uncaptioned image] The Faroe Islands are an archipelago in the North Atlantic Ocean and an autonomous territory of the Kingdom of Denmark. Located between Iceland, Norway, and the United Kingdom, the islands have a population of 54,900 as … Instance of (P31) Country (Q6256)
Country (P17) Denmark (Q35)
Part of (P361) Northern Europe (Q27479)
Continent (P30) Europe (Q46)
Capital (P36) Tórshavn (Q10704)
Stephen
Curry
[Uncaptioned image] Wardell Stephen Curry II, born March 14, 1988, also known as Steph Curry, is an American professional basketball player and point guard for the Golden State Warriors of the National Basketball Association (NBA). Instance of (P31) Human (Q5)
Sex/Gender (P21) Male (Q6581097)
Citizenship (P27) United States (Q30)
Occupation (P106) Basketball P. (Q3665646)
Father (P22) Dell Curry (Q1184381)
Lady
Gaga
[Uncaptioned image] Stefani Joanne Angelina Germanotta[a] (born March 28, 1986), known professionally as Lady Gaga, is an American singer, songwriter, and actress. She is an influential figure in popular music, known for her image reinventions and … Instance of (P31) Human (Q5)
Sex/Gender (P21) Female (Q6581072)
Citizenship (P27) United States (Q30)
Occupation (P106) Singer (Q177220)
Occupation (P106) Songwriter (Q753110)
Milan
Kundera
[Uncaptioned image] Milan Kundera was a Czech and French novelist. Kundera went into exile in France in 1975, acquiring citizenship in 1981. His Czechoslovak citizenship was revoked in 1979, but he was granted Czech citizenship in 2019… Instance of (P31) Human (Q5)
Sex/Gender (P21) Male (Q6581097)
Citizenship (P27) France (Q142)
Occupation (P106) Writer (Q36180)
Instrument (P1303) Piano (Q5994)
Qualitative Analysis
Input Image Text Query Prediction Ground Truth
1 2 3
what is this building called?
Q6406516
Q15278727
Q7637054
Q6406516
which species of insect is this?
Q5135458
Q5379231
Q4694221
Q4694221
where is this place?
Q1014274
Q2193788
Q1012308
Q2075156
what is the model of this aircraft?
Q6423
Q62161
Q852
Q6423
what is the name of this park?
Q517545
Q912770
Q7381109
Q517545
Table 8: Example KnowCoL’s predictions on five test images. For each query, the model’s top-3 (according to the cosine similarity) Wikidata items predictions are shown (columns 1–3), with green borders indicating correct matches and red borders indicating incorrect ones; the ground-truth item is in the rightmost column.
To complement our quantitative evaluation, we present qualitative examples of KnowCoL’s predictions on the OVEN test set in Table 8. These cases illustrate the model’s ability to correctly recognize fine-grained entities under zero-shot settings.
Overall, KnowCoL can place the correct Wikidata Q-ID at rank 1 for most cases, even though those three are drawn from a pool of more than 30000 candidates. Its main failure occurs when the top candidates are similar (e.g., different medieval ruins), causing the true entity to slip out of the top three entirely. We also observe that the top three candidates are all highly plausible alternatives to the ground truth.